# AI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions

### v1

@@ 8 Discussion
We highlighted the importance of deepfakes in the field of digital forensics, which we consider to be a topic of great relevance and timeliness: in particular, the concept of impostor bias that we introduced is increasingly prevalent, although it had not yet been named and defined precisely. A concrete example of impostor bias was seen with the recent war in Ukraine, the first war characterized by constant propaganda from both sides carried out on all media, and especially on social networks (Ciuriak, 2022; Suciu, 2022). Being flooded with videos, photos and statements inevitably leads to believe a priori that everything that is seen can be modified, or generated by AI models through a simple prompt (Linehan et al., 2023). The habit of suspicion developed during this “digital warfare” with propaganda and false information led to immediate doubt of all published media, sometimes leading to the discovery that some of this media was deepfake (Bond, 2023).
This condition of prior doubt on the veracity of the media will be more and more invasive and constant, especially when it comes to delicate material, such as a communication from a head of state in wartime (Linehan et al., 2023): for this reason, we believe that it is necessary to use the best tools for detecting deepfakes and AI-generated multimedia to avoid being deceived, especially to avoid considering fake multimedia that, paradoxically, are instead real.
Therefore, the data presented herein, along with previous findings, underscore the imperative need for further exploration of cognitive biases in digital forensic science. The focus should not be solely on the biases themselves, but rather on the strategies that can be employed to counteract and disengage from them. One potential approach could involve the creation and implementation of specialized training programs for forensic practitioners. These programs would aim to address cognitive biases in various scenarios, such as those examined in this review. By enhancing awareness of cognitive biases, we could potentially reduce the likelihood of practitioners being influenced or overwhelmed by these biases. This would enable them to disregard irrelevant elements of the investigation, separate personal experiences and other contextual factors that may “contaminate” decision-making processes, thereby preserving the objectivity of their judgments.
@@ 9 Conclusions and Future Works
This article explores cognitive biases in forensics and digital forensics, including confirmation bias, expectation bias, and overconfidence in errors. It discusses methods to reduce these biases, such as game-based interventions and the Linear Sequential Unmasking-Expanded approach. The article highlights the challenge of deepfakes, synthetic media that can manipulate or impersonate individuals, threatening the integrity of digital evidence. It surveys current deepfake detection methods and their limitations. The article introduces the impostor bias, a new cognitive bias that may lead to false negatives and reduced confidence in digital forensic findings due to the prevalence of deepfakes. We conclude by proposing new research directions in digital forensics, suggesting the creation of advanced deepfake detection methods, exploring the factors contributing to impostor bias, and recommending interventions to mitigate this bias. We also underscore the need to examine the ethical, legal, and social implications of deepfakes.

### v2

@@ 8 Discussion
The concept of Impostor Bias, though not yet widely recognized, is increasingly relevant in the era of deepfakes and digital forensics. The recent Ukraine war, marked by propaganda and misinformation on social media (Ciuriak, 2022; Suciu, 2022), has heightened this bias. The constant exposure to videos, photos, and statements fosters an inherent suspicion of media manipulation or AI generation (Linehan et al., 2023). This ”digital warfare” has led to immediate doubt and scrutiny of all published media, sometimes uncovering deepfakes (Bond, 2023). The prevalence of Impostor Bias will persist, especially with sensitive content like wartime communications (Linehan et al., 2023). Thus, novel techniques and strategies are essential for effective mitigation.
The unique nature of the Impostor Bias necessitates specialized approaches beyond the strategies discussed in Section 4. While those strategies offer valuable insights, addressing Impostor Bias effectively requires a combination of targeted measures. This includes shielding analysts from irrelevant information and providing specific training on cognitive bias risks. Additionally, proficiency testing and regular assessments maintain practitioner competence. However, to truly counter Impostor Bias, advanced technical tools for synthetic content detection are essential. Practical deep learning algorithms can detect generative models’ traces, aiding investigators in distinguishing authentic evidence. Given the success of deepfake detection algorithms (Pei et al., 2024; Gong and Li, 2024), these tools can assist in counteracting Impostor Bias and other digital biases. Similar detection methods for forgery images (Baumy et al., 2022; Zanardelli et al., 2023; Singh and Kumar, 2024) and multimedia data manipulation (Galante et al., 2023; Dunsin et al., 2024) further reinforce the value of algorithmic assistance in avoiding bias. The increasing sophistication of digital data demands the use of advanced deepfake detection techniques to maintain trust and accuracy in forensic examinations.
Detection methods play a crucial role in identifying deepfakes, but interpreting and explaining the results is equally important. While some methods use classic machine learning techniques, others employ deep learning techniques, each with its advantages and limitations. Classic machine learning approaches are more interpretable but lack robustness, while deep learning methods offer higher performance and robustness but are more complex to explain. The question of trusting the predictions of detection algorithms is indeed a separate but related study, as explored by Mathews et al. (2023), Lim et al. (2022), and others (Cantero-Arjona and Sánchez-Macián, 2024; Pinhasov et al., 2024; Pandey et al., 2024; Pontorno et al., 2024). These studies enhance the “trust” in detection methods and underscore the need for further exploration of cognitive biases in digital forensic science. Creating specialized training programs for practitioners can enhance awareness, reduce the influence of biases, and preserve the objectivity of judgments by disregarding irrelevant elements and separating personal experiences.
@@ 9 Conclusions and Future Works
In this article, we have explored the significant impact of cognitive biases in the fields of forensics and digital forensics, specifically focusing on confirmation bias, anchoring bias, and hindsight bias. To reduce the influence of these biases, we have discussed strategies such as game-based interventions and the Linear Sequential Unmasking-Expanded approach. With the growing prevalence of deepfakes, synthetic media that can manipulate or impersonate individuals, the integrity of digital evidence is at risk. We have surveyed current deepfake detection methods and their limitations, introducing the novel concept of Impostor Bias. This bias, influenced by the widespread use of deepfakes, may lead to false negatives and reduced confidence in digital forensic findings. To address these challenges, we propose future research directions, including the development of advanced deepfake detection methods and a deeper understanding of the factors contributing to Impostor Bias. Furthermore, we emphasize the importance of interventions to mitigate this bias and the need to explore the ethical, legal, and social implications of deepfakes. By doing so, we aim to enhance the reliability and integrity of digital forensic practices, ensuring the accuracy and objectivity of forensic investigations.